OpenStudy (anonymous):

log(4x-1)=log(x+1)+log2

OpenStudy (anonymous):

the right hand side is the same as \(\log(2(x+1))\) so your actual job is to solve \[4x-1=2(x+1)\] for \(x\)

OpenStudy (anonymous):

thanx!